Systems Staff Meeting Notes
January 24, 2005
January 24, 2005
Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Geoffrey StewartAnnouncements
Ongoing Systems Updates
- Apocalyptic discussion of global warming/global dimming. The earth receives 22% less sunlight than in the 1950's.
- Brief ALEPH outage over the weekend. Bill placed a call to FCLA and it was expeditiously fixed.
- Bobbie visited the Butterfly Rainforest over the weekend. Her granddaughter was reprimanded for touching a dead butterfly.
- AC - has been out in room 16 since late Thursday afternoon. PPD just got here to repair.
- Back ordered printers for DLC - both have been received, as well as 20 Lacie hard drives. All were purchased on a PO. They have not been checked in yet due to a lingering question of procedure.
- Greenstone order should go out this week - institutional repository project (slated to be renamed). It will be 8-terrabyte, expandable, Linux, Australian software - probably an IBM server.
- Spyware - big response from liaisons - Alexa and DSO exploit are the most common. Until we can render the spyware catchers centrally updateable, we'll let the liaisons continue to monitor and clean machines. We don't want to end up with the troubles we had with virus checkers. Will wait until we can push it from our end.
Software Projects Status
- Network gremlin still not found. Bill is suspicious of network hardware (switches). Have been experiencing painfully slow File Transfer. CLAS is also having troubles. SQL server access seems good at the moment.
- NT4 changeover - NT 4.0 server converted to Win 2003. Will took a laptop and joined the domain. It connected fine. He took it to the North Tower and it still worked. Think we can start moving things to the new domain. Primary domain controller has to be moved first. This should eliminate the use of WINS. Will have to re-point individual workstations to internal DNS server. Moving management over to CNS should stabilize. MS is holding a seminar on converting to NT4 next month. There will be major conversions this year. Some things will break. Will get rid of LAN management at the same time.
- ALEPH printing - Will received an email from Carol Turner - they are planning training groups and want Will's input on the technical aspects of ALEPH printing. Seems to be a form of group therapy - everyone is able to print, some notices are not correct, but that is not a technical issue. As far as we're concerned, it's working as advertised.
- Computers in West - should be getting feedback from the committee soon.
- ALEPH/PeopleSoft permissions - dismal conclusion to meeting last Friday. Still trying to straighten out permission problem with CIRC person in Education - no listings in his Dean's Menu. Bill called 4-5 people, had to have the user call PeopleSoft directly.
- Jack Waters - data transfer problem. Transactions were recorded, but there is still a problem on the Bridges end.
- Discussion on PeopleSoft employees losing their jobs with the Oracle takeover.
- Position Description Database - is finished. Winston will set it up to grab all of the permissions from the Personnel database. Positions must be defined on paper.
- DLC - nothing major.
Student and Department Happenings
- Upcoming: new outreach committee established by Carol Turner focusing on communicating with users on the Web. Push for video involvement. Bill Hanssen, Administration's graphic designer, will be assisting in that regard.
- Barbara, Sandy and Dale would all like a splashy, new page to introduce the new Library West. In using video, you must have a written transcript. Tom will probably use Flash.
- All is well. Seem to all be in action this week.